how to find a unknown resistance by post box method

Grade:12

1 Answers

Vikas TU
6 years ago
It is a Wheatstone Bridge with three arms P, Q and R; while the fourth arm(s) is the obscure resistance. P and Q are known as the proportion arms while R is known as the rheostat arm. 
At adjust, the obscure resistance 
S = (P/Q) R … (1) 
The proportion arms are first balanced with the goal that they convey 100 W each. The resistance in the rheostat arm is currently balanced so that the galvanometer diversion is in one heading. In the event that R = R0 (ohm) and 
R = R0 + 1 (ohm) are the resistance in rheostat arm, for which the diversion in galvanometer is in inverse bearing, then it suggests that the obscure resistance "s" lies between R0 and (R0 + 1) ohm. 
Alt txt: resistance-utilizing post-office-box 
Presently, the resistance in P and Q are made 100 W and 1000 W, individually, and the procedure is rehashed. Condition (1) is utilized to register S. The proportion P/Q is dynamically made 1:10, and afterward 1:100.